# Nonlinear Causality in Brain Networks: With Application to Motor Imagery vs Execution

### v1

@@ Abstract
One fundamental challenge of data-driven analysis in neuroscience is modeling causal interactions and exploring the connectivity of nodes in a brain network. To this end, many statistical methods, relying on various perspectives and employing different data modalities, are being developed to examine and comprehend the underlying causal structures inherent to brain dynamics. This study develops a novel statistical approach enabling neuroscientists to dissect causal interactions in multichannel EEG recordings. Our approach, TAR4C, uses the threshold autoregressive model to describe the causal interaction between nodes or clusters of nodes in a brain network. The perspective involves testing whether one node, which may represent a brain region, can control the dynamics of the other. The node that has such an impact on the other is called a threshold variable and can be classified as a causative because its functionality is the leading source operating as an instantaneous switching mechanism that regulates the time-varying autoregressive structure of the other. This statistical concept is commonly referred to as threshold non-linearity. Once threshold non-linearity has been verified between a pair of nodes, the subsequent essential facet of TAR modeling is to assess the predictive ability of the causal node for the current activity on the other and represent causal interactions in autoregressive terms. This predictive ability is what underlies Granger causality. The TAR4C approach can discover non-linear and time-dependent causal interactions without negating the G-causality perspective. The efficacy of the proposed approach is exemplified by analyzing the EEG signals recorded during the motor movement/imagery experiment. The similarities and differences between the causal interactions manifesting during the execution and the imagery of a given motor movement are demonstrated by analyzing EEG recordings from multiple subjects.
@@ 1 Introduction
The exploration and understanding of the complex connectivity and causality intrinsic to networks of time series data represents a pivotal challenge that necessitates significant methodological enhancements, particularly within the domains of neuroscience and statistical analyses for brain imaging data (Hamedi et al., 2016; Sheikhattar et al., 2018; Farahani et al., 2019; Runge et al., 2019; Biswas and Shlizerman, 2022; Varley and Sporns, 2022). As our understanding of brain function deepens, it becomes increasingly evident that brain connectivity is both dynamic and nonlinear, demanding sophisticated analytical approaches to decode the underlying interactions (Razi and Friston, 2016; Cao et al., 2022; Siddiqi et al., 2022). The further progress and analytical capabilities attained for this objective are also propelling progress in other fields, including economics and climate science, when and where the objective is to analyze time series data from a network/graph perspective (Papana et al., 2017; Gong et al., 2019; Upadhyay et al., 2020; Wang and Huang, 2021; Nowack et al., 2020; Naghipour et al., 2021; Holme and Rocha, 2023; Tjøstheim et al., 2023).
This study aims to propose a conceptually comprehensive methodology that captures nonlinear causal connectivities between any given Nodes of Interest (NOIs) within a network of time series. To align with this aim, a data analytical objective is set to explore causal interactions within a network of electroencephalogram (EEG) time series data. While conventional and state-of-the-art methods are undoubtedly practical and helpful, they are limited in their ability to incorporate the nonlinear, especially dynamic phase-shifting (i.e., phase transition) and time-varying nature of these interactions (Jirsa et al., 1994; Beggs and Plenz, 2003; Beggs, 2008, 2022; Corsi et al., 2024; Shriki et al., 2013; Sporns, 2016; Bocaccio and Tagliazucchi, 2024).
@@ 1.1 Literature Background Specific to Causal Inference in EEG
A significant number of methodologies, including Granger Causality (GC) (Ding et al., 2006; Marinazzo et al., 2011; Barrett et al., 2012; Zhao et al., 2013; Seth et al., 2015; Benozzo et al., 2017; Stokes and Purdon, 2017; Sheikhattar et al., 2018; Dhamala et al., 2018; Zhang et al., 2019; Wismüller et al., 2021; Heyse et al., 2021; Biswas and Ombao, 2022), Transfer Entropy (TE) (Vicente et al., 2011; Huang et al., 2015; Diks and Fang, 2017; Ursino et al., 2020; Redondo et al., 2023), Directed Transfer Function (DTF) (Babiloni et al., 2005; Wilke et al., 2008; Wang et al., 2016; Heyse et al., 2021; Li and Zhang, 2022), Partial Directed Coherence (PDC) (Chiang et al., 2009; Varotto et al., 2012; Omidvarnia et al., 2013; He et al., 2014; Mill et al., 2017; Gaxiola-Tirado et al., 2017), Dynamic Causal Modeling (DCM) (Friston et al., 2003; Kiebel et al., 2008; Chen et al., 2008; Daunizeau et al., 2009; Chicharro and Panzeri, 2014; Van de Steen et al., 2019; Wang and Liang, 2021), Bayesian modeling (Prado et al., 2001; Rajapakse and Zhou, 2007; Gorrostieta et al., 2012, 2013; Yu et al., 2016; Chiang et al., 2017; Kook et al., 2020; Hu et al., 2020; Granados-Garcia et al., 2024), and Machine / Deep Learning (ML / DL) modeling (Schirrmeister et al., 2017; Chai et al., 2019; Phang et al., 2019; Zhang et al., 2019, 2021; Gao et al., 2021; Rosol et al., 2022; Luo et al., 2022; Liu et al., 2023; Bocaccio and Tagliazucchi, 2024) based approaches, have been proposed as means of inferring causality and connectivity in EEG, with varying degrees of focus on specific objective aspects.
Furthermore, approach-based literature surveys outlined by Haufe et al. (2011); Razi and Friston (2016); Hamedi et al. (2016); Farahani et al. (2019); Ombao and Pinto (2022); Siddiqi et al. (2022); Shojaie and Fox (2022); Biswas and Shlizerman (2022); Chiarion et al. (2023); Zheng et al. (2023) offer valuable insights into the current framework on human brain connectivity.
@@
The conventional GC method employs linear autoregressive models, which assess whether the past values of one time series can enhance the prediction of another. It should be noted that due to the inherent linearity and Gaussianity assumption of conventional GC, its usage in the complex, nonlinear dynamics of brain signals will lead to limited applications and interpretations (Lusch et al., 2016; Shojaie and Fox, 2022). In light of these structural restrictions, researchers have sought to develop alternative approaches that would permit the G causality framework to be extended to the nonlinear dynamics characterizing real-world contexts (e.g., Marinazzo et al., 2011; Zhao et al., 2013; Henderson and Michailidis, 2014; Chen et al., 2017; Runge et al., 2019; Wismüller et al., 2021; Tank et al., 2022). Developing methodology for Nonlinear GC inference remains an active area of research due to several inherent challenges. These include the difficulty of determining the degree of nonlinearity, the prevalence of overfitting issues, the incompatibility of the method with high dimensionality, the generation of high variance estimators, the necessity for a large number of hyperparameters, the use of black-box routines, and the complexity of interpreting the results.
@@
On the other hand, TE is a nonparametric, model-free method for quantifying directional dependencies between time series, particularly useful for capturing nonlinear interactions. However, TE faces challenges related to computational complexity, data requirements, and interpretability. It requires large datasets for reliable estimates, and its sensitivity to noise and data sparsity, especially in EEG, can lead to false positives or spurious connections. TE’s performance is further influenced by parameter choices, such as embedding delay and dimension, adding complexity. In multivariate systems, TE may mistakenly attribute indirect connections as direct, leading to inaccurate conclusions. Partial Transfer Entropy (PTE) addresses this issue by considering other variables but retains many of TE’s computational and interpretative challenges. Despite these limitations, TE’s ability to reveal hidden causal relationships makes it a powerful tool across disciplines, though ongoing refinement is needed to improve its practicality and clarity (Wibral et al., 2014; Ekhlasi et al., 2023; Das and Porfiri, 2023).
Dynamic Causal Modeling (DCM) is a Bayesian framework used in neuroimaging to model how brain regions influence each other based on neural and hemodynamic principles. Unlike Granger causality, DCM uses mechanistic differential equations to describe neural activity over time and integrates prior knowledge about brain structure to estimate causal influences. While these biologically grounded priors allow for detailed inferences about brain dynamics, they also risk model misspecification if the assumptions are incorrect, leading to biased results. Additionally, DCM’s computational intensity and sensitivity to the selection of priors make model comparison challenging, as different priors can lead to varying conclusions. Despite these limitations, DCM is valued for its ability to incorporate biophysical constraints, making it particularly useful for testing specific hypotheses about neural circuits, such as the directionality of connections during cognitive tasks (Weichwald and Peters, 2021; Lohmann et al., 2012; Razi and Friston, 2016).
@@
In contrast, machine learning models have gained popularity for their ability to detect complex, nonlinear relationships in EEG data, but they are often criticized for their lack of interpretability. These models can effectively classify and predict brain states by recognizing subtle patterns in high-dimensional data. However, they often operate as ”black boxes,” offering little insight into the underlying causal mechanisms of neural activity. This ambiguity can impede the extraction of meaningful insights about causal relationships in brain networks. Moreover, machine-learning approaches require large, well-labeled datasets, which are often scarce in neuroscience. EEG data are particularly noisy, and machine learning models are highly sensitive to this noise, which can lead to poor generalization across different subjects, experimental conditions, or datasets if the overfitting issue is not properly addressed. Furthermore, while ML methods are adept at capturing associations and providing strong predictive power, distinguishing between causality and mere correlation remains elusive, requiring the integration of additional causal reasoning frameworks. Reproducibility across studies can also be a concern, as model assumptions, dataset variability, and preprocessing differences can result in inconsistent findings across studies (Wein et al., 2021; Ji et al., 2021). Despite these challenges, ML’s flexibility and capacity for handling vast amounts of data offer exciting prospects, especially when combined with rigorous causal inference methods.
@@ 1.2 Challenges of Inferring Causal Interactions in EEG Data
Inferring causal interactions in brain activity is challenging due to the brain’s complex, nonlinear, and time-varying dynamics. Connectivity constantly changes with stimuli, cognitive processes, and internal conditions, making it difficult to separate direct causal relationships from correlations and interdependence.
It is well-documented that brain functioning is characterized by various types of nonlinearities (Mayer-Kress, 1998; Breakspear and Terry, 2002; Breakspear, 2017; Lynn and Bassett, 2019; O’Donnell, 2023). Neural nonlinearities are complex, non-proportional relationships in brain function that appear in various forms (e.g., synaptic plasticity, neural oscillations, neuronal firing rates). The literature on the exploration of brain connectivity has sought to encompass the phenomenon of nonlinearity through the use of state-of-the-art approaches and empirical investigations. However, the notion that the brain functions in a non-linear way is taken as an umbrella concept, and it is assumed that each study addresses nonlinearity appropriately. Since nonlinearity can take various forms, it may be more accurate to refer to ’nonlinearities’ rather than ’nonlinearity’ in brain function. A single methodology cannot capture all these variations or support unified conclusions. This paper analyzes the abrupt phase-shifting dynamics of the brain from an unexplored nonlinearity perspective in relation to EEG data, aligning with findings that the brain processes information with nonlinear, abrupt phase shifts (Beggs, 2008; Razi and Friston, 2016).
@@
A further challenge in inferring causality from EEG data resides in the limited capacity of EEG to depict spatial resolution of neuronal activities in sufficient detail. This insufficiency makes it difficult to localize neural sources, complicating the task of identifying genuine causal relationships between brain regions. The recorded signals often reflect aggregated activity from multiple sources, distorted by volume conduction, which can create undecomposable correlations and spurious causal inferences, especially for nearby channels. While methods like source localization and beamforming can help mitigate these issues, they also introduce their own limitations and assumptions that must be carefully considered. The causal inference obtained from EEG often reflects a simplified model of brain connectivity, lacking the granularity needed to understand neural mechanisms comprehensively. While EEG reveals the directional flow of information between channels, it offers limited insight into the biophysical processes involved, such as synaptic or neurotransmitter activity. Therefore, EEG-based causality should be interpreted cautiously, as it represents statistical connections rather than direct neural interactions. Despite its significant temporal resolution, EEG’s limited spatial and mechanistic detail restricts its value in brain connectivity studies. However, it remains useful for exploring neural timing and sequence, especially when combined with other methods (Bastos and Schoffelen, 2016).
1.3 Objective of the Study
This study aims to introduce a new approach, the Threshold Autoregressive model for Causality (TAR4C), to discover nonlinear and time-varying causal interactions, especially when the goal is to dissect the connectivity between only a few nodes in multi-node networks of time series, focusing on EEG data. The specific objectives are as follows:
Introducing Threshold Nonlinearity for Connectivity Inference: Existing methods overlook a key brain principle: abrupt transition dynamics, often focusing on specific aspects of nonlinearity. In the TAR framework, the threshold variable (i.e., node/channel) triggers activity in another node when its nonlinearity effect is significant. The threshold variable’s activity governs the other, suggesting a connection between the two nodes, even without additional evidence. This approach reveals phase transitions through threshold variables, offering a new understanding of EEG connectivity and dynamics.
@@ 1.3 Objective of the Study
Modeling Nonlinear and Time-Varying Causal Interactions: Another goal is to model how different nodes in a network influence each other under different conditions by validating nonlinear and time-varying interactions. Applying TAR4C to EEG data enables us to study the causal relationships as threshold variable’s states change, particularly during motor tasks like execution and imagery. TAR4C is, therefore, capable of tracking causal connectivity alterations over time, which is not addressed by current methodologies.
@@
Comparing Connectivity Patterns in Motor Execution and Imagery: The study also aims to examine differences in brain connectivity during motor execution (MM) and motor imagery (MI). It investigates how causal interactions between regions, such as the prefrontal and motor cortices, vary across tasks involving different gestures (e.g., left fist, right fist, both fists, both feet). By applying TAR4C, the study discovers distinct patterns in how the brain responds to real and imagined motor tasks.
@@
Quantifying Connectivity with New Indices: The TAR4C method introduces two new metrics: the Threshold Connectivity Index (TCI) and the Threshold Granger Causality Index (TGCI). These indices measure the strength and direction of causal interactions between EEG channels (e.g., nodes). The TCI identifies threshold-based connectivity, while the TGCI evaluates the predictive relationships between channels. These indices provide a simplified representation of nonlinear brain interactions.
@@
Expanding TAR4C Beyond EEG Networks: Although focused on EEG data, the TAR4C framework can be applied to other time series networks, like those in economics or climate, when the goal is to uncover relationships between specific variables in high-dimensional settings. This study positions TAR4C as a general tool for detecting nonlinear causal interactions across various fields, emphasizing the broader relevance of threshold autoregressive models in understanding complex connectivity patterns.
2 Threshold Autoregressive Modeling for Exploring Connectivity and Causality
@@
This study proposes an approach based on threshold autoregressive (TAR) modeling to explore connectivity within time series networks and a novel application to multi-channel motor EEG recordings presented as a case study. The brain, a highly dynamic structure, constantly shifts between different connectivity states in response to stimuli and tasks. EEG devices measure brain activity through signals captured by electrodes (channels/nodes) on the scalp. These signals, rich in temporal information, are valuable data for studying brain connectivity during cognitive processes, neurological disorders, and responses to various stimuli.
EEG recordings reflect time-varying connectivities and phase transitions that occur during brain activities, such as perception, cognition, and motor tasks. This study assumes that these transitions are governed by phase-shifting mechanisms that form effective connections between brain regions. The TAR model, designed to handle regime shifts, can track these transitions by specifying functionals that control the dynamics. By identifying these phase changes, the TAR model provides a practical tool for discovering causal interactions among EEG channels, offering insights into how different channels influence one another.

### v2

@@ Abstract
One fundamental challenge of data-driven analysis in neuroscience is modeling causal interactions and exploring the connectivity between nodes in a brain network. Various statistical methods, using different perspectives and data modalities, have been developed to understand the causal structures in brain dynamics. This study introduces a novel statistical approach, TAR4C, to dissect causal interactions in multichannel EEG recordings. TAR4C uses the threshold autoregressive (TAR) model to describe causal interactions between nodes in a brain network from two perspectives. The first tests whether one node controls the dynamics of another. The controlling node, named the threshold variable, implies its causative role since it operates as a switching mechanism governing the instantaneous transitions between autoregressive structures. This concept is known as threshold non-linearity. Once verified between a node pair, the next step in TAR modeling is assessing the causal node’s predictive ability on the other’s activity, representing causal interactions in autoregressive terms, a concept underlying Granger (G) causality. TAR4C can discover non-linear, time-dependent causal interactions while maintaining the G-causality framework. The approach’s efficacy is demonstrated through EEG data from a motor execution/imagery experiment. By comparing causal interactions during motor execution and imagery, TAR4C reveals key similarities and differences in brain connectivity across subjects.
@@ 1 Introduction
The exploration of the complex connectivity and causality intrinsic to networks of time series data represents a pivotal challenge that necessitates significant methodological enhancements, particularly within the domain of neuroscience (Biswas and Shlizerman, 2022). Many experiments over the last few decades have produced increasing evidence that brain connectivity is both dynamic and nonlinear, thus demanding sophisticated analytical approaches to decode the underlying interactions. This study is motivated by the challenging problem of differentiating brain networks between motor imagery vs. motor execution tasks. This has a potential impact on identifying the role of brain function (in particular, connectivity dynamics) in the treatment of stroke. Our aim is to propose a conceptually comprehensive methodology that captures non-linear causal connectivity between any given nodes of interest (NOIs) within a network of time series. To align with this aim, a key analytical objective is set to explore (Granger) causal interactions within a network of electroencephalogram (EEG) time series data. While conventional and state-of-the-art methods are undoubtedly practical and helpful, they are limited in their ability to incorporate the nonlinear, especially dynamic phase-shifting (i.e., phase transition) and time-varying nature of these interactions (Sporns, 2016).
@@ 1.1 Literature Background Specific to Causal Inference in EEG
A significant number of methods have been proposed as means of inferring causality and connectivity in EEG, with varying degrees of focus on specific objective aspects, including Granger Causality (GC)(Seth et al., 2015; Ombao and Pinto, 2022), Transfer Entropy (TE) (Vicente et al., 2011; Redondo et al., 2023), Directed Transfer Function (DTF) (Babiloni et al., 2005), Partial Directed Coherence (PDC) (He et al., 2014), Bayesian modeling (Prado et al., 2001; Gorrostieta et al., 2012; Yu et al., 2016; Chiang et al., 2017), and Machine / Deep Learning (ML / DL) modeling (Shojaie and Fox, 2022)
@@
The conventional Granger Causality method employs linear autoregressive models, which assess whether the past values of one time series can improve the prediction of another (as determined by lessening in squared prediction error). Due to the inherent linearity assumption of conventional GC, its usage in the complex, nonlinear dynamics of brain signals will lead to limited applications and interpretations (Shojaie and Fox, 2022). To overcome these restrictions, alternative approaches have been developed to permit the Granger causality framework to be extended to the nonlinear dynamics characterizing real-world contexts (e.g., Tank et al., 2022). The challenges to developing methodology for nonlinear GC include the difficulty of determining the degree of nonlinearity, the prevalence of overfitting issues, the incompatibility of the method with high dimensionality, the generation of high variance estimators, the necessity for a large number of hyperparameters, the use of black-box routines, and the complexity of interpreting the results.
@@
On the other hand, TE is a nonparametric, model-free method for quantifying directional dependencies between time series, particularly useful for capturing nonlinear interactions. However, TE faces challenges related to computational complexity, data requirements, and interpretability. It requires large datasets for reliable estimates, and its sensitivity to noise and data sparsity, especially in EEG, can lead to false positives or spurious connections.
@@
In contrast, machine learning models have gained popularity for their ability to detect complex, nonlinear relationships in EEG data, but they are often criticized for their lack of interpretability. These models can effectively classify and predict brain states but they often operate as ”black boxes,” offering little insight into the underlying causal mechanisms of neural activity. Moreover, machine-learning approaches require large, well-labeled datasets, which are often scarce in neuroscience. Reproducibility across studies can also be a concern, as model assumptions, dataset variability, and preprocessing differences can result in inconsistent findings across studies (Wein et al., 2021). Despite these challenges, ML’s flexibility and capacity for handling vast amounts of data offer exciting prospects, especially when combined with rigorous causal inference methods.
@@ 1.2 Challenges of Inferring Causal Interactions in EEG Data
Inferring causal interactions in brain activity, as it responds to stimuli, poses several challenges, largely due to the brain’s complex, nonlinear, and time-varying dynamics. Brain functioning is characterized by various types of nonlinearities (Breakspear, 2017) Neural nonlinearities are complex relationships in brain function that appear in various forms (e.g., synaptic plasticity, neural oscillations, neuronal firing rates). This paper aims to analyze the abrupt phase-shifting dynamic of the human brain from a previously unexplored and undiscussed nonlinearity perspective in relation to EEG data. The approach we propose is in line with the findings that the brain operates with nonlinearity characterized by abrupt phase shifts when processing information.
A further challenge in inferring causality from EEG data resides in the limited capacity of EEG to depict spatial resolution of neuronal activities in sufficient detail. Thus, neural sources are not localized complicating the task of identifying genuine causal relationships between brain regions. The recorded signals often reflect aggregated activity from multiple sources, distorted by volume conduction, which can create undecomposable correlations and spurious causal inferences.
@@
Source localization and beamforming can help mitigate these issues but they also introduce their own limitations and assumptions that must be carefully considered. Moreover, the causal inference derived from EEG often reflects a simplified model of brain connectivity. While EEG reveals the directional flow of information between channels, it offers limited insight into the biophysical processes involved, such as synaptic or neurotransmitter activity. Therefore, the causal relationships inferred from EEG data should be interpreted with caution, as they represent statistical associations rather than direct neural interactions (Bastos and Schoffelen, 2016; Ombao and Pinto, 2022).
1.3 Objective of this Paper
We develop a new approach, the Threshold Autoregressive model for Causality (TAR4C), to discover nonlinear and time-varying causal interactions in a high dimensional network. The TAR4C method minds the gap in existing methods, which can only comprehend several manifestations of nonlinearities. By using a threshold nonlinearity perspective, TAR4C offers a way to explore brain dynamics, particularly phase transitions in brain connectivity, which reflect how a reference channel may impact the interaction between nodes. The specific objectives are as follows:
Introducing Threshold Nonlinearity for Connectivity Inference. The existing methods fail to consider the operating principle of the brain, which is characterized by abrupt transition dynamics, and tend to prioritize particular facets of nonlinearity. The threshold variable (i.e., node/channel) in the TAR framework serves as a triggering reference for the other node. In other words, the level of activity in the threshold variable governs the other variable’s activity. It can thus be reasoned that the two nodes (e.g., regions) are connected in this respect alone. This approach identifies phase transitions based on threshold (e.g., cause) variables, providing a new way to understand the brain’s connectivity and dynamic processes.
@@ 1.3 Objective of this Paper
Modeling Nonlinear and Time-Varying Causal Interactions. Another goal is to model how different nodes in a network influence each other under different conditions by validating nonlinear and time-varying interactions. Applying TAR4C to EEG data enables us to study and test the causal relationships as threshold variable’s states change, particularly during motor tasks like execution and imagery. TAR4C is, therefore, capable of tracking causal connectivity alterations over time, which is not addressed by current methodologies.
@@
Comparing Connectivity Patterns in Motor Execution and Imagery. The study also aims to examine differences in brain connectivity during motor execution (ME) and motor imagery (MI). It investigates how causal interactions between regions, such as the prefrontal and motor cortices, vary across tasks involving different gestures (e.g., open and close of both fists, both feet). By applying TAR4C, the study discovers distinct patterns in how the brain responds to real and imagined motor tasks.
@@
Quantifying Connectivity with New Indices. The TAR4C method introduces two new metrics: the Threshold Connectivity Index (TCI) and the Threshold Granger Causality Index (TGCI). These indices measure the strength and direction of causal interactions between EEG channels (e.g., nodes). The TCI identifies threshold-based connectivity, while the TGCI evaluates the predictive relationships between channels. These indices provide a simplified representation of nonlinear brain interactions.
@@
Expanding TAR4C Beyond EEG Networks. Although the data analytic focus is on EEG data, the TAR4C framework can also be applied to other networks of time series, such as those pertaining to economics or climate variables, especially when the objective is to discover relationships between specific variables in a high-dimensional setting. This study establishes TAR4C as a tool for detecting nonlinear causal interactions across various domains, highlighting the broader relevance of threshold autoregressive models in understanding complex connectivity patterns.
2 Exploring Connectivity and Causality via Threshold Autoregressive Models
@@
This study proposes an approach based on threshold autoregressive (TAR) modeling to explore connectivity within time series networks and a novel application to multi-channel motor EEG recordings presented as a case study. EEG devices measure brain activity through signals captured by electrodes (channels/nodes) on the scalp. These signals, rich in temporal information, are valuable data for studying brain connectivity during tasks (e.g., motor). The brain, a highly dynamic structure, constantly shifts between different connectivity states in response to stimuli and tasks. This study assumes that phase transitions are governed by phase-shifting mechanisms that form effective connections between brain regions. The TAR model, designed to handle regime shifts, can track these transitions by specifying functionals that control the dynamics. By identifying these phase changes, the TAR model provides a practical tool for discovering causal interactions among EEG channels, offering insights into how different channels influence one another.